# Are LLMs Good Literature Review Writers? Evaluating the Literature Review Writing Ability of Large Language Models

### v2

@@ 4.1 Experimental Settings
LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3 and Llama-3.2-3B-Instruct.
@@ 4.2 Main Results
Results for Task 1. By analyzing the results in Table 1, we observe that Claude-3.5 achieves the highest accuracy and title search rate, whereas Llama-3.2 exhibits the lowest performance on both metrics for Task 1. Furthermore, when evaluating the accuracy of LLM-generated references, treating the author item as correct when the first author is identified correctly results in a 3–5% increase in accuracy scores across all models. This finding underscores that accurately generating complete author lists remains a significant challenge for LLMs. Additionally, we conduct a year-wise analysis of the literature that each model cited correctly, as illustrated in the line graph in Figure 2. The graph shows that the majority of the correctly cited references are concentrated between 2010 and 2020.
Results for Task 2. We use various metrics to evaluate model performance. As shown in Table 2, Claude-3.5 consistently achieves the best overall performance across nearly all metrics. The average semantic similarity between the abstracts generated by Claude-3.5 and human-written abstracts is 81.17%. Furthermore, Claude-3.5 demonstrates the highest factual consistency, with a TURE model evaluation score of 78.10% and a GPT-4o assessment score of 96.77%. Additionally, we observe that Llama-3.2 achieves the highest score on the ROUGE-L metric, but does not show a clear advantage on other metrics. Given the diversity of outputs produced by LLMs, it is crucial to employ multiple evaluation metrics when assessing their performance on our tasks.
Results for Task 3. As shown in Table 3, we observe that, compared to the accuracy scores in Task 1, all LLMs exhibit a significant increase in accuracy when generating references in Task 3. Previous studies have shown that citing real external sources can reduce hallucination rates in the generated text. In our experiment, we find that when LLMs are tasked with generating references for the literature review, the accuracy of the generated references increases significantly. This suggests that the generated references and the literature review text impose constraints on each other.
On the other hand, we find that Claude-3.5 achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5 recalls the most claims from the human-written literature review. While the literature review generated by DeepSeek-V3 performs well on the ROUGE metrics.
@@
Table 1: The experimental results of the four LLMs in Task 1. “$S_{t}$” refers to the title search rate as defined in Eq 5, while “Acc.” represents the accuracy defined in Eq 4. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
@@ 4.3 Analyze LLM-Generated References from Different Dimensions
In both Task 1 and Task 3, we ask LLMs to generate references. The overall performance was discussed in the previous section. In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2. As seen in Figure LABEL:f2a, Claude-3.5 demonstrates a clear advantage over other models across all dimensions in Task 1. Additionally, the accuracy of reference generation in Task 1 for Claude-3.5, GPT-4, and Qwen-2.5 follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension. Accuracy for journal name, page, and author is also relatively high. However, DeepSeek-V3 performs worse in the author dimension compared to the other dimensions. In contrast, Llama-3.2 demonstrates higher accuracy in the page and author dimensions than in other dimensions. However, overall, Llama-3.2 does not exhibit a competitive advantage in reference generation accuracy.
Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5, DeepSeek-V3, GPT-4o, and Qwen-2.5, with particularly obvious gains in the author dimension. The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension. Notably, the accuracy of DeepSeek-V3 and GPT-4 in certain dimensions approaches or even exceeds that of Claude-3.5. However, the performance of LLaMA-3.2 remains suboptimal.
@@ 4.4 Cross-Disciplinary Analysis
First, we observe that in Task 1, as shown in Figure LABEL:f4a, almost all models exhibit the highest accuracy in the Mathematics discipline and lowest accuracy in the Chemistry discipline. Additionally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences. The results are as follows: Claude-3.5 (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5 (p<.0001), Llama-3.2 (p=0.065). These results indicate significant differences in performance across disciplines for each model except Llama-3.2.
Secondly, as shown in Figure LABEL:f4b, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2. We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology. Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines. The results are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5 (p<.001), Llama-3.2 (p<.0001). These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
Additionally, we examine the accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c. It is evident that the accuracy of Claude-3.5, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5 and LLaMA-3.2 across all disciplines. Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology. The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5(p<.0001), Llama-3.2(p<.001), indicating significant differences in each models’ performance across the various disciplines.
Finally, we observe the KPR metric across different disciplines in Task 3, as shown in Figure LABEL:f4d. The results from the figure indicate that the differences between models—Claude-3.5, DeepSeek-V3, GPT-4, and Qwen-2.5—are not significant across various disciplines, a finding that is also supported by statistical tests. Specifically, the p-values are as follows: Claude-3.5 (p=0.46), DeepSeek-V3 (p=0.23), GPT-4 (p=0.18), Qwen-2.5 (p=0.10), and Llama-3.2 (p<0.001).
@@ 4.5 Comparison of LLM-Cited and Human-cited References
In this section, we also compare the LLM-generated references with those from the human-written original articles. Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article. As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5 are also cited in the human-written article, indicating a relatively high overlap rate. Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3. Notably, GPT-4o models show a 10% increase in the overlap rate. Next, we provide a more detailed comparison of the references across various dimensions. As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower. For Task 3, Claude-3.5 and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1. This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
@@ Appendix B Data Distribution
Figure 4: Statistics of dataset.

### v3

@@ 4.1 Experimental Settings
LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3, and Llama-3.2-3B-Instruct.
@@ 4.2 Main Results
Results for Task 1. By analyzing the results in Table 1, we observe that Claude-3.5-Sonnet achieves the highest accuracy and title search rate, whereas Llama-3.2-3B exhibits the lowest performance on both metrics for Task 1. Furthermore, when evaluating the accuracy of LLM-generated references, treating the author item as correct when the first author is identified correctly results in a 3–5% increase in accuracy scores across all models. This finding underscores that accurately generating complete author lists remains a significant challenge for LLMs. Additionally, we conduct a year-wise analysis of the literature that each model cited correctly, as illustrated in the line graph in Figure 2. The graph shows that the majority of the correctly cited references for almost all LLMs involved in the task are concentrated between 2010 and 2020.
Results for Task 2. We use various metrics to evaluate model performance. As shown in Table 2, Claude-3.5-Sonnet consistently achieves the best overall performance across nearly all metrics. The average semantic similarity between the abstracts generated by Claude-3.5-Sonnet and human-written abstracts is 81.17%. Furthermore, Claude-3.5-Sonnet and DeepSeek-V3 exhibit the highest factual consistency, achieving a TURE model evaluation score of 78.10% and a GPT-4o assessment score of 96.84%, respectively. Additionally, we observe that Llama-3.2-3B achieves the highest score on the ROUGE-L metric but does not exhibit a clear advantage on other metrics. This suggests that, given the diversity of outputs produced by LLMs, it is crucial to employ multiple evaluation metrics when assessing their performance on our tasks.
Results for Task 3. As shown in Table 3, we observe that, compared to the accuracy scores in Task 1, all LLMs exhibit a significant increase in accuracy when generating references in Task 3. Previous studies have shown that citing real external sources can reduce hallucination rates in the generated text (Gao et al., 2023). In our experiment, we find that when LLMs are tasked with generating references for the literature review, the accuracy of the generated references increases significantly. This suggests that the generated references and the literature review text impose constraints on each other.
On the other hand, we find that Claude-3.5-Sonnet achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5-Sonnet recalls the most claims from the human-written literature review. While the literature review generated by DeepSeek-V3 performs well on the ROUGE metrics.
@@
Table 1: The experimental results of the five LLMs in Task 1. “$S_{t}$” refers to the title search rate as defined in Eq 4, while “Acc.” represents the accuracy defined in Eq 3. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
@@ 4.3 Analyze LLM-Generated References from Different Dimensions
In both Task 1 and Task 3, we ask LLMs to generate references. The overall performance was discussed in the previous section. In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2. As seen in Figure LABEL:f2a, Claude-3.5-Sonnet demonstrates a clear advantage over other models across all dimensions in Task 1. Additionally, the accuracy of reference generation in Task 1 for Claude-3.5-Sonnet, GPT-4o, and Qwen-2.5-72B follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension. Accuracy for journal name, page, and author is also relatively high. However, DeepSeek-V3 performs worse in the author dimension compared to the other dimensions. In contrast, Llama-3.2 demonstrates higher accuracy in the page and author dimensions than in other dimensions. However, overall, Llama-3.2-3B does not exhibit a competitive advantage in reference generation accuracy.
Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B, with particularly obvious gains in the author dimension. The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension. Notably, the accuracy of DeepSeek-V3 and GPT-4o in certain dimensions approaches or even exceeds that of Claude-3.5-Sonnet. However, the performance of LLaMA-3.2-3B remains suboptimal.
@@ 4.4 Cross-Disciplinary Analysis
First, we observe that in Task 1, as shown in Figure LABEL:f4a, almost all models exhibit the highest accuracy in the Mathematics discipline and lowest accuracy in the Chemistry discipline. Additionally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences. The results are as follows: Claude-3.5-Sonnet (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B (p<.0001), Llama-3.2-3B (p=0.065). These results indicate significant differences in performance across disciplines for each model except Llama-3.2-3B.
Secondly, as shown in Figure LABEL:f4b, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2. We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology. Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines. The results are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5-72B (p<.001), Llama-3.2-3B (p<.0001). These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
Additionally, we examine the references accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c. It is evident that the accuracy of Claude-3.5-Sonnet, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5-72B and LLaMA-3.2-3B across all disciplines. Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology. The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B(p<.0001), Llama-3.2-3B(p<.001), indicating significant differences in each models’ performance across the various disciplines.
Finally, we observe the KPR metric across different disciplines in Task 3, as shown in Figure LABEL:f4d. The results from the figure indicate that the differences between models—Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B—are not significant across various disciplines, a finding that is also supported by statistical tests. Specifically, the p-values are as follows: Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p<0.001).
@@ 4.5 Comparison of LLM-Cited and Human-cited References
In this section, we also compare the LLM-generated references with those from the human-written original articles. Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article. As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5-Sonnet are also cited in the human-written article, indicating a relatively high overlap rate. Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3. Notably, GPT-4o models show a 10% increase in the overlap rate. We provide a more detailed comparison of the references across various dimensions in Appendix C.
4.6 Human Evaluation
To evaluate the reliability of our automatic assessment method for identifying hallucinated references, we conduct a comparative analysis involving 100 LLM-generated references. These references were assessed by three annotators and the final manual results were obtained by majority vote. The results demonstrated a kappa agreement of 0.71 between the automatic and human assessments, signifying a relatively high level of consistency and supporting the reliability of our method. Furthermore, when using human assessment results as the gold standard, the automatic assessment method achieved an accuracy of 86%, further validating its effectiveness.
@@ Appendix B Data Distribution
Statistics of the dataset are shown in Figure 4.
Appendix C Comparison of LLM-Cited and Human-cited References from Different Dimensions
We provide a more detailed comparison of the LLM-cited and human-cited references across various dimensions. As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower. For Task 3, Claude-3.5-Sonnet and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1. This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
Appendix D Discussion